# Linear and Sublinear Diversities

### v2

@@ 2.3 Characterization of linear diversities
The following characterization of linear diversities is essentially contained in the proof of the main theorem in Firey [10]; see also [21].
@@
Let $\delta$ be a function defined on finite subsets of $\mathbb{R}^{k}$. Then $(X,\delta)$ is a linear semidiversity if and only if there is a positive finite Borel measure $\nu$ on the unit sphere $\mathbb{S}^{k-1}=\{x\in\mathbb{R}^{k}:\|x\|_{2}=1\}$ such that
@@
First we show that $\delta$ given by (4) is a linear semidiversity. For $a\in\mathbb{R}^{k}$ and finite $A\subseteq B\subseteq\mathbb{R}^{k}$ we have
$\delta(\{a\})$ $=\int_{\mathbb{S}^{k-1}}h_{\{a\}}(x)\,\mathrm{d}\nu(x)=\int_{\mathbb{S}^{k-1}}a\cdot x\,\mathrm{d}\nu(x)=a\cdot\int_{\mathbb{S}^{k-1}}x\,\mathrm{d}\nu(x)=0$
and, since $h_{A}(x)\leq h_{B}(x)$ and $h_{A+B}(x)=h_{A}(x)+h_{B}(x)$ for all $x$ we have $\delta(A)\leq\delta(B)$ and $\delta(A+B)=\delta(A)+\delta(B)$. By Proposition 1, $(\mathbb{R}^{k},\delta)$ is a linear semidiversity.
For the converse, let $(\mathbb{R}^{k},\delta)$ be a linear semidiversity and define $\delta^{*}$ as in (2). By Proposition 4 the restriction of $\delta^{*}$ to nonempty compact convex subsets is Minkowski linear, monotonic and vanishes on singletons. From the proof of the main theorem in [10], we have that, for all compact convex sets $K$,
@@
for some positive finite Borel measure $\nu$ satisfying (3), and $\nu$ is the unique such measure. (See [23, Thm 2.14] for details on the use of the Riesz Theorem in this case.) Now for any nonempty finite $A$, let $K=\mathrm{conv}(A)$. Since $\delta(A)=\delta^{*}(K)$ and $h_{A}=h_{K}$, the result follows for all nonempty finite $A$. ∎
Figure 1: Support of measures corresponding to (a) mean width; (b) the $L_{1}$ diversity; and (c) a Minkowski diversity with a simplex kernel.
In Figure 1 we depict the support for the measures corresponding to mean width (uniform on the unit circle), the $L_{1}$ diversity (±$\pm e_{i}$), and the Minkowski diversity for a simplex kernel. The first two of these are easy enough to demonstrate. We prove the third example below, after we have a characterization of extremal linear diversities.
2.4 Extremal linear diversities
The set of linear semidiversities on $\mathbb{R}^{k}$ forms a cone. A non-zero semidiversity $\delta$ is extremal (or lies on an extremal ray) if it cannot be expressed as the convex combination of two linear semidiversities which are not its scale copies. We make use of Theorem 5 to characterize the extremal linear diversities and semidiversities. First we prove a technical result simplifying evaluation of the Minkowski diversity for a simplex.
Lemma 6.
Let $v_{0},\ldots,v_{j}\in\mathbb{R}^{k}$ be affinely independent with $\sum_{\ell=0}^{j}c_{\ell}v_{\ell}=0$, for some $c_{\ell}\geq 0$, $\sum_{\ell}c_{\ell}=1$. Define the polyhedron $K=\{y:v_{\ell}\cdot y\leq 1,\,\text{for all }\ell\}$. Let $\delta_{K}$ be the Minkowski semidiversity given by $K$. Then
$\delta_{K}(A)=\sum_{\ell=0}^{j}c_{\ell}h_{A}(v_{\ell}).$
for all finite $A\subseteq\mathbb{R}^{k}$.
@@ 2.4 Extremal linear diversities
Let $A=\{a_{i}\}_{i=1,\ldots,|A|}$. We express $\delta_{K}(A)$ as the solution to a linear program. Recall that $\delta_{K}(A)$ is the minimum $\lambda$ such that there is some $x\in\mathbb{R}^{k}$ such that $a_{i}-x\in\lambda K$ for all $i$. We can rewrite this constraint as $v_{\ell}\cdot(a_{i}-x)\leq\lambda$ for all $i,\ell$. If we take $\lambda$ and $x$ to be our primal variables we get the following linear program:
minimize $\lambda=(1,0)\cdot(\lambda,x)$
subject to $\lambda+v_{\ell}\cdot x\geq v_{\ell}\cdot a_{i},\mbox{for all }i\mbox{ and }\ell.$
The dual linear program with dual variables $y_{i\ell}$ is
maximize $\sum_{i\ell}(v_{\ell}\cdot a_{i})y_{ij}$
subject to $y_{i\ell}\geq 0,\mbox{for all }i\mbox{ and }\ell,$
$\sum_{i\ell}y_{i\ell}=1,$
$\sum_{i\ell}v_{\ell}y_{i\ell}=0.$
Let $\bar{y}_{\ell}=\sum_{i}y_{i\ell}$. Then our dual constraints are equivalent to
$\sum_{\ell}\bar{y}_{\ell}=1,\ \ \ \sum_{\ell}\bar{y}_{\ell}v_{\ell}=0.$
Since the $v_{\ell}$ are affinely independent and $\sum_{\ell}c_{\ell}v_{\ell}=0$, $\sum_{\ell}c_{\ell}=1$, there is a unique solution given by $\bar{y}_{\ell}=c_{\ell}$ for all $\ell$. Now it remains to determine for each $\ell$ the value of $y_{i\ell}$ for each $i$. We need to maximize $\sum_{i\ell}(v_{\ell}\cdot a_{i})y_{i\ell}$ given $y_{i\ell}\geq 0$ and $\sum_{i}y_{i\ell}=c_{\ell}$. For each $\ell$, the solution is to let $y_{i\ell}=c_{\ell}$ for the $i$ that maximizes $(v_{\ell}\cdot a_{i})$, and 00 otherwise. This gives for the solution to the dual problem
$\sum_{\ell}c_{\ell}\max_{a\in A}v_{\ell}\cdot a=\sum_{\ell}c_{\ell}h_{A}(v_{\ell}).$
The following theorem identifies extremal linear semidiversities as Minkowski diversities $\delta_{K}$ with $K$ equal to a simplex or a simplex plus a subspace.
Theorem 7.
The following are equivalent for a semidiversity $(\mathbb{R}^{k},\delta)$:
  1. (i)
    $(\mathbb{R}^{k},\delta)$ is extremal in the class of linear semidiversities.
  2. (ii)
    $(\mathbb{R}^{k},\delta)$ satisfies
    $\delta(A)=\int_{\mathbb{S}^{k-1}}h_{A}(x)\,\mathrm{d}\nu(x)$
    for all finite $A\subseteq\mathbb{R}^{k}$, where $\nu$ is a measure on $\mathbb{S}^{k-1}$ with $\int_{\mathbb{S}^{k-1}}x\,\mathrm{d}\nu(x)=0$, such that the support of $\nu$ is a finite, affinely independent set.
  3. (iii)
    $(\mathbb{R}^{k},\delta)$ is a Minkowski semidiversity with kernel $K$ of the form
    $K=\mathrm{conv}(W)+H^{\perp},$
    where $W$ is an affinely independent set of points, $H$ is the affine closure of $W$,and $H^{\perp}$ is the orthogonal space to $H$.
Proof.
(i) $\Rightarrow$ (ii). Suppose $\delta$ is extremal and the support of $\nu$ is not affinely independent. Let $H$ be the affine hull of the support of $\nu$, with $\dim H=j$, and let $S=H\cap\mathbb{S}^{k-1}$. Affine dependence implies $\nu$ is not supported on only $j+1$ points or fewer. Therefore we can partition $S$ into $S_{1},\ldots,S_{j+2}$, each with $\nu(S_{i})>0$. Let $m_{i}=\int_{S_{i}}\,\mathrm{d}\nu(x)=\nu(S_{i})>0$. Then
[…]
and
$0=\int_{S}x\,\mathrm{d}\nu(x)=\sum_{j}\int_{S_{i}}x\,\mathrm{d}\nu(x)=\sum_{i}m_{i}x_{i}$
where $x_{i}=(\int_{S_{i}}x\,\mathrm{d}\nu(x))/m_{i}$. Choose a subset of the $x_{i}$ with $k+1$ points so that 0 is in the convex hull of them. Let’s say they are $x_{1},\ldots,x_{k+1}$. Find $\mu_{i}$ for $i=1,...,k+1$ such that $\sum_{i}\mu_{i}x_{i}=0$, and $\mu_{i}<m_{i}$. Let $\mu_{k+2}=0$. Now define $\nu^{\prime}$ by $\nu^{\prime}(A)=(\mu_{j}/m_{j})\nu(A)$ for $A\subseteq S_{i},j=1...k+1$, and zero otherwise. Then $\nu^{\prime}\leq\nu$, and $\nu^{\prime}$ has smaller support, because $m_{k+2}>0$ but $\mu_{k+2}=0$. Also
[…]
We can now write $\nu=(\nu-\nu^{\prime})+\nu^{\prime}$ where $\nu-\nu^{\prime}$ and $\nu$ are not scale copies, so $\delta$ is not extremal in the cone of linear semidiversities.
(ii) $\Rightarrow$ (i). Suppose that
$\delta(A)=\int_{\mathbb{S}^{k-1}}h_{A}(x)\,\mathrm{d}\nu(x)$
for all finite $A\subseteq\mathbb{R}^{k}$ and some measure $\nu$ on $\mathbb{S}^{k-1}$ with affinely independent support. Let $\delta_{1}$ and $\delta_{2}$ be linear semidiversities with corresponding measures $\nu_{1}$ and $\nu_{2}$. If $\mathrm{supp}(\nu_{1})\subseteq\mathrm{supp}(\nu)$ then affine independence and the constraint that $\int_{\mathbb{S}^{k-1}}x\,\mathrm{d}\nu_{1}(x)=0$ implies that $\nu_{1}$ is a scaled version of $\nu$. Likewise for $\nu_{2}$. Hence if $\nu=\lambda\nu_{1}+(1-\lambda)\nu_{2}$ for $\lambda\in[0,1]$ we have that $\mathrm{supp}(\nu_{1})\cup\mathrm{supp}(\nu_{2})\subseteq\mathrm{supp}(\nu)$ and both $\nu_{1}$ and $\nu_{2}$ are scale versions of $\nu$. This shows that $\delta$ is an extremal linear diversity.
(ii) $\Rightarrow$ (iii). Let the support of $\nu$ be the points $u_{0},\ldots,u_{j}$ with weights $m_{\ell}>0$ such that $\sum_{\ell}m_{\ell}u_{\ell}=0$. Let $m=\sum_{\ell}m_{\ell}$, $c_{\ell}=m_{\ell}/m$, and $v_{\ell}=mu_{\ell}$, so that $\delta(A)=\sum_{\ell}c_{\ell}h_{A}(v_{\ell})$, $\sum_{\ell}c_{\ell}v_{\ell}=0$, and $\sum_{\ell}c_{\ell}=1$. Let $V=\{v_{0},\ldots,v_{j}\}$, which is affinely independent because the $u_{\ell}$ are. Let $H$ be the span of $V$ and $H^{\perp}$ its orthogonal complement. By Lemma 6, $\delta=\delta_{K}$, the Minkowski semidiversity for the set $K=\{y:y\cdot v_{\ell}\leq 1,\mbox{ for all }\ell\}$. The intersection of $K$ with $H$ is a simplex; let it have vertices $W=\{w_{0},\ldots,w_{j}\}$. Then $K=\mathrm{conv}(V)+H^{\perp}$ as required.
(iii) $\Rightarrow$ (ii). By translating $K$ if necessary, we may assume 00 is in the relative interior of $\mathrm{conv}(W)$. We can write $K=\{y:v_{\ell}\cdot y\leq 1\}$ for some affinely independent $V=\{v_{0},\ldots,v_{j}\}$. Because $\mathrm{conv}(W)$ is bounded, $0\in\mathrm{conv}(V)$. So there are $c_{\ell}\geq 0$ with $\sum_{\ell}c_{\ell}v_{\ell}=0$ and $\sum_{\ell}c_{\ell}=1$. By Lemma 6 we have
$\delta_{K}(A)=\sum_{\ell}c_{\ell}h_{A}(v_{\ell})$
for all finite $A$. Let $m_{\ell}=c_{\ell}|v_{\ell}|$ and $u_{\ell}=v_{\ell}/|v_{\ell}|$. Then the $u_{\ell}$ are also affinely independent. Let $\nu$ be the measure that assigns mass $m_{\ell}$ to each $u_{\ell}$. ∎
Points in a finite dimensional convex cone can always be written as convex combinations of extremal points. The cone of linear semidiversities has infinite dimensional, so proving that linear semidiversities are in the convex hull of extremal diversities requires a little more work.
Theorem 8.
A semidiversity $(\mathbb{R}^{k},\delta)$ is linear if and only if $\delta$ is a convex combination of extremal linear semidiversity functions.
Proof.
Since a weighted average of linear semidiversities is a linear semidiversity, one way is immediate. For the other, suppose that $\delta$ is a linear semidiversity. By Theorem 5, there is a Borel measure $\nu$ on $\mathbb{S}^{k-1}$ such that $\int x\,\mathrm{d}\nu(x)=0$ and $\delta(A)=\int h_{A}(x)\,\mathrm{d}\nu(x)=0$ for all finite $A$. Let $m=\int\,\mathrm{d}\nu(x)$.
Let $E$ be the set of all signed Borel measures on $\mathbb{S}^{k-1}$, which is a Hausdorff locally convex set [26, p. 134]. The space $C$ of measures on $\mathbb{S}^{k-1}$ with $\int x\,\mathrm{d}\nu(x)=m$ and $\int x\,\mathrm{d}\nu(x)=0$ is compact by the Banach-Alaoglu theorem [26, p. 114], and is convex.
We claim that the set of extremal points of $C$ is closed. Let $\nu_{n}$, $n\geq 1$ be a sequence of extremal measures that converges in the vague topology, so that $\int f\,\mathrm{d}\nu_{n}$ converges to $\int f\,\mathrm{d}\nu$ for some $\nu\in C$ for all continuous bounded $f$. By repeatedly taking subsequences, we can obtain a subsequence $\nu_{n_{k}}=\sum_{i=1}^{j}\mu_{i,k}\delta_{x_{i,k}}$ (where $\delta_{x}$ is a unit mass measure at $x$) where $x_{i,k}\rightarrow x_{i}$ and $\mu_{i,k}\rightarrow\mu_{i}$ for some $x_{i}\in\mathbb{S}^{k-1}$ and $\mu_{i}\geq 0$. Since $\int f\,\mathrm{d}\nu_{n_{k}}\rightarrow\int f\,\mathrm{d}\nu$ as $k\rightarrow\infty$, we must have $\nu=\sum_{i}\mu_{i}\delta_{x_{i}}$, showing that the limit is also an extremal point in $C$. Hence the set of extremal measures is closed.
We can apply a version of the Krein-Milman Theorem ([26, Corollary 17.7]) to obtain that $\delta$ is a weighted average of members of the closure of the extreme points of $C$. Since the set of extremal points of $C$ is closed, the result follows. ∎
@@ 2.5 Characterization of sublinear diversities
We now turn our attention to sublinear diversities. We will show that the relationship between sublinear and linear diversities parallels that between convex and linear functions. Just as every convex function is the supremum of linear functions, every sublinear diversity is the supremum of linear diversities (Theorem 9). In fact, in our case, the supremum is attained for each set, so the value of every sublinear diversity on a set is the maximum of the value of a family of linear diversities on the set. Our proof relies heavily on the ‘Sandwich Theorem’ (Theorem 1.2.5) of [11].
@@
Conversely, for any collection $\{(\mathbb{R}^{k},\delta_{\gamma})\}_{\gamma\in\Gamma}$ of linear semidiversities and $\delta$ defined by $\delta(A)=\sup_{\gamma\in\Gamma}\delta_{\gamma}(A)$, $(\mathbb{R}^{k},\delta)$ is a sublinear semidiversity.
@@
$\delta(A+B)=\sup\{\delta_{\gamma}(A+B):\gamma\in\Gamma\}=\sup\{\delta_{\gamma}(A)+\delta_{\gamma}(B):\gamma\in\Gamma\}\leq\delta(A)+\delta(B).$
So $\delta$ is sublinear. By Proposition 1, $\delta$ is a sublinear semidiversity.
For the converse, suppose that $(\mathbb{R}^{k},\delta)$ is sublinear. Define $\mathcal{H}$ to be the set of all support functions $h_{A}$ for nonempty finite $A\subseteq\mathbb{R}^{k}$. Define $p$ on the convex cone $\mathcal{H}$ by $p(h_{A})=\delta(A)$ for all finite sets $A$. The function $p$ is sublinear (and convex in the terminology of [25]), as for any finite $A,B$,
@@
and $p(\lambda h_{A})=\lambda p(h_{A})$ for $\lambda\geq 0$.
@@
$q_{B}(h_{A})=\sup\{\lambda:\lambda B+x\subseteq\mathrm{conv}(A)\mbox{ for some$x\in\mathbb{R}^{k}$}\}.$
That is, $q_{B}(h_{A})$ is the largest we can scale $B$ so that a translate is contained in $\mathrm{conv}(A)$. Note that $q_{B}(h_{B})=1$. This tells us that $p(h_{B})=\delta(B)=\delta(B)q_{B}(h_{B})$.
We show that $q_{B}$ is superlinear. For all $\alpha\geq 0$ we have $q_{B}(\alpha h_{A})=q_{B}(h_{\alpha A})=\alpha q_{B}(h_{A})$. Now suppose that $A_{1},A_{2}$ are finite and non-empty subsets of $\mathbb{R}^{k}$. Given $\epsilon>0$ there are $\lambda_{1}>q_{B}(h_{A_{1}})-\epsilon/2$, $\lambda_{2}>q_{B}(h_{A_{2}})-\epsilon/2$, $x_{1},x_{2}\in\mathbb{R}^{k}$ such that
@@
We now have that $p$ is monotonic and sublinear and that $q_{B}$ is superlinear.
Furthermore, for any finite $A\subseteq\mathbb{R}^{k}$ and $\epsilon>0$ there is $\lambda$ such that $q_{B}(h_{A})-\epsilon<\lambda\leq q_{B}(h_{A})$ and $x\in\mathbb{R}^{k}$ such that $\lambda B+x\subseteq\mathrm{conv}(A)$, and so
$p(h_{A})$ $\geq p(h_{\lambda B})$
$=\lambda p(h_{B})$
$>(q_{B}(h_{A})-\epsilon)\delta(B).$
@@
For each finite $B$ we have now satisfied the conditions for Theorem 1.2.5 of [11]:
@@
$\delta(B)q_{B}(h)\leq\mu_{B}(h)\leq p(h).$
Since by definition $p(h_{\{a\}})=\delta(\{a\})=0$ for all $a\in\mathbb{R}^{k}$, $\mu_{B}(h_{\{a\}})=0$ for all $a\in\mathbb{R}^{k}$.
Now define $\delta_{B}$ by $\delta_{B}(A)=\mu_{B}(h_{A})$ for all finite $A$. Then $\delta_{B}$ vanishes on singletons, it is monotone, linear, and hence also sublinear. By Proposition 1 $(\mathbb{R}^{k},\delta_{B})$ is a linear semidiversity.
@@
$\delta_{B}(B)=\mu_{B}(h_{B})=\delta(B)q_{B}(B)=\delta(B)$

### v3

@@ 2.3 Characterization of linear diversities
The following characterization of linear diversities is essentially contained in the proof of the main theorem in Firey firey1976functional . A clear statement of the relevant version of the Riesz representation theorem may be found in rotem2021riesz ; see also meyer1995convex ; hug2020lectures .
@@
Let $\delta$ be a function defined on finite subsets of $\mathbb{R}^{k}$. Then $(\mathbb{R}^{k},\delta)$ is a linear semidiversity if and only if there is a positive finite Borel measure $\nu$ on the unit sphere $\mathbb{S}^{k-1}=\{x\in\mathbb{R}^{k}:\|x\|_{2}=1\}$ such that
@@
First we show that $\delta$ given by (8) is a linear semidiversity. For $a\in\mathbb{R}^{k}$
$\delta(\{a\})$ $=\int_{\mathbb{S}^{k-1}}h_{\{a\}}(x)\,\mathrm{d}\nu(x)=\int_{\mathbb{S}^{k-1}}a\cdot x\,\mathrm{d}\nu(x)=a\cdot\int_{\mathbb{S}^{k-1}}x\,\mathrm{d}\nu(x)=0.$
For finite $A,B\subseteq\mathbb{R}^{k}$ and $\lambda\geq 0$, $h_{\lambda A}=\lambda h_{A}$ and $h_{A+B}=h_{A}+h_{B}$ so $\delta(\lambda A)=\lambda\delta(A)$ and $\delta(A+B)=\delta(A)+\delta(B)$. If $A\subseteq B$ then $h_{A}(x)\leq h_{B}(x)$ for all $x\in\mathbb{R}^{k}$, giving $\delta(A)\leq\delta(B)$ since $\nu$ is positive. By Proposition 1, part 2., $(\mathbb{R}^{k},\delta)$ is a linear semidiversity.
For the converse, let $(\mathbb{R}^{k},\delta)$ be a linear semidiversity and define $\delta^{*}$ as in (6). By Proposition 3 the restriction of $\delta^{*}$ to nonempty compact convex subsets is Minkowski linear and monotonic. The Lipschitz property in Proposition 3 part 5 means we can use (rotem2021riesz, , Thm. 1.2) to obtain that there is a unique positive finite Borel measure $\nu$ such that
@@
for all compact convex sets $K$. Note that for all $a\in\mathbb{R}^{k}$ we have
$0=\delta^{*}(\{a\})=\int_{\mathbb{S}^{k-1}}a\cdot x\,\mathrm{d}\nu(x)=a\cdot\int_{\mathbb{S}^{k-1}}\ x\,\mathrm{d}\nu(x)$
and so $\int_{\mathbb{S}^{k-1}}\ x\,\mathrm{d}\nu(x)=0$. Now for any nonempty finite $A$, let $K=\mathrm{conv}(A)$. Then $\delta(A)=\delta^{*}(K)$ by Proposition 3, part 2., and $h_{K}=h_{A}$, so
$\delta(A)=\delta^{*}(K)=\int_{\mathbb{S}^{k-1}}h_{K}(x)\,\mathrm{d}\nu(x)=\int_{\mathbb{S}^{k-1}}h_{A}(x)\,\mathrm{d}\nu(x),$
as required. ∎
Figure 2: Support of measures corresponding to (a) mean width; (b) the $\ell_{1}$ diversity; and (c) a Minkowski diversity with a simplex kernel.
In Figure 2 we depict the support for the measures corresponding to three different linear diversities:
  1. (a)
    The mean width diversity, where $\nu$ is the uniform Haar measure on the sphere.
  2. (b)
    The $\ell_{1}$ diversity, where $nu$ has point masses at $\pm e_{i}$, since
    $\delta(A)$ $=\int_{\mathbb{S}^{k-1}}h_{A}(x)\,\mathrm{d}\nu(x)=\sum_{i=1}^{k}(h_{A}(e_{i})+h_{A}(-e_{i}))=\sum_{i=1}^{k}\max_{a,b\in A}|a_{i}-b_{i}|.$
  3. (c)
    The minimal case, in the sense that the support of $\nu$ is a set of affinely independent points, equivalently, the vertices of a simplex. No linear diversity has a smaller support. The corresponding diversities are characterized in the following Proposition.
Proposition 6.
Let $v_{0},\ldots,v_{k}\in\mathbb{R}^{k}$ be affinely independent and suppose that $\sum_{i=0}^{k}c_{i}v_{i}=0$ for positive $c_{0},\ldots,c_{k}$ such that $\sum_{i=0}^{k}c_{i}=1$. Define the polyhedron $K=\bigcap_{i=0}^{k}\{y\in\mathbb{R}^{k}:y\cdot v_{i}\leq 1\}$. Let $\delta_{K}$ be the Minkowski diversity with kernel $K$. Then
$\delta_{K}(A)=\sum_{i=0}^{k}c_{i}h_{A}(v_{i})$
for all finite $A\subseteq\mathbb{R}^{k}$. It follows that $(\mathbb{R}^{k},\delta_{K})$ is a linear diversity.
@@
Suppose that $A=\{a_{1},a_{2},\ldots,a_{m}\}$. We express $\delta_{K}(A)$ as the solution to a linear program which is a special case of a more general formulation in BrandenbergRoth09 :
$\text{min}_{\lambda,x}\quad$ $\lambda$
subject to $\lambda+v_{i}\cdot x\geq v_{i}\cdot a_{j}$
 $\quad\mbox{ for all }i=0,1,\ldots,k\mbox{ and }j=1,\ldots,m.$
The dual linear program has variables $y_{ij}$ for $i=0,1,\ldots,k$ and $j=1,\ldots,m$:
$\text{max}_{y}\quad$ $\sum_{i=0}^{k}\sum_{j=1}^{m}(v_{i}\cdot a_{j})y_{ij}$
subject to $y_{ij}\geq 0,\mbox{ for all $i=0,\ldots,k$ and $j=1,\ldots,m$}$
$\sum_{i=0}^{k}\sum_{j=1}^{m}y_{ij}=1,$
$\sum_{i=0}^{k}\sum_{j=1}^{m}v_{i}y_{ij}=0.$
Define $z_{i}=\sum_{j=1}^{m}y_{ij}\geq 0$ for all $i=0,\ldots,k$. Then our dual constraints are equivalent to
$\sum_{i=0}^{k}z_{i}=1,\ \ \ \sum_{i=0}^{k}z_{i}v_{i}=0.$
Since the $v_{\ell}$ are affinely independent and $\sum_{i=0}^{k}c_{i}v_{i}=0$, $\sum_{i=0}^{k}c_{i}=1$ we have $z_{i}=c_{i}$ for all $i=0,\ldots,k$ when $y$ is feasible.
For each $i=0,\ldots,k$ the maximum of $\sum_{j=1}^{m}(v_{i}\cdot a_{j})y_{ij}$ such that $y_{i1},\ldots,y_{im}$ are non-negative and $z_{i}=\sum_{j=1}^{m}y_{ij}=c_{i}$ is obtained when $y_{ij^{*}}=c_{i}$ for $j^{*}$ maximizing $(v_{i}\cdot a_{j})$ and $y_{ij}=0$ for $j\neq j^{*}$. This maximization can be carried out independently for each $i=0,\ldots,k$. Hence the optimal value for the dual and primal problems is
$\sum_{i=0}^{k}c_{i}\max_{a_{j}\in A}v_{i}\cdot a_{j}=\sum_{i=0}^{k}c_{i}h_{A}(v_{i}),$
as required. ∎
@@ 2.4 Characterization of sublinear diversities
We now turn our attention to sublinear diversities. We will show that the relationship between sublinear and linear diversities parallels that between convex and affine functions. Just as every convex function is the supremum of affine functions, every sublinear diversity is the supremum of linear diversities (Theorem 7). In fact, in our case, the supremum is attained for each set, so the value of every sublinear diversity on a set is the maximum of the value of a family of linear diversities on the set. Our proof relies heavily on the ‘Sandwich Theorem’ (Theorem 1.2.5) of FuchssteinerLusky81convex .
@@
Conversely, for any collection $\{(\mathbb{R}^{k},\delta_{\gamma})\}_{\gamma\in\Gamma}$ of linear semidiversities and $\delta$ defined by $\delta(A)=\sup_{\gamma\in\Gamma}\delta_{\gamma}(A)$, the pair $(\mathbb{R}^{k},\delta)$ is a sublinear semidiversity.
@@
$\delta(A+B)=\sup\{\delta_{\gamma}(A+B):\gamma\in\Gamma\}=\sup\{\delta_{\gamma}(A)+\delta_{\gamma}(B):\gamma\in\Gamma\}\leq\delta(A)+\delta(B),$
so $(\mathbb{R}^{k},\delta)$ is sublinear. By Proposition 1, $(\mathbb{R}^{k},\delta)$ is a sublinear semidiversity.
For the converse, suppose that $(\mathbb{R}^{k},\delta)$ is sublinear. Define $\mathcal{H}$ to be the set of all support functions $h_{A}$ for nonempty finite $A\subseteq\mathbb{R}^{k}$. Define $p$ on the convex cone $\mathcal{H}$ by $p(h_{A})=\delta(A)$ for all finite sets $A$. By Proposition 1 part 3, $\delta(A)$ only depends on $\mathrm{conv}(A)$ so $p$ is well-defined on $\mathcal{H}$. The function $p$ is sublinear as for any finite $A,B$,
@@
and $p(\lambda h_{A})=p(h_{\lambda A})=\delta(\lambda A)=\lambda\delta(A)=\lambda p(h_{A})$ for $\lambda\geq 0$.
@@
$q_{B}(h_{A})=\sup\{\lambda:\lambda B+x\subseteq\mathrm{conv}(A)\mbox{ for some $x\in\mathbb{R}^{k}$}\}$
with $q_{B}(h_{A})=0$ if $|A|=1$. When $|A|\geq 2$, $q_{B}(h_{A})$ is the largest we can scale $B$ so that a translate is contained in $\mathrm{conv}(A)$. Note that $q_{B}(h_{B})=1$ and hence $p(h_{B})=\delta(B)=\delta(B)q_{B}(h_{B})$.
We show that $q_{B}$ is superlinear. For all $\alpha\geq 0$ we have $q_{B}(\alpha h_{A})=q_{B}(h_{\alpha A})=\alpha q_{B}(h_{A})$.
Let $A_{1},A_{2}$ be finite, non-empty subsets of $\mathbb{R}^{k}$. If $|A_{1}|=1$ then $A_{1}+A_{2}$ is a translate of $A_{1}$, so
$q_{B}(A_{1}+A_{2})=q_{B}(A_{2})=q_{B}(A_{1})+q_{B}(A_{2}).$
Likewise, if $|A_{2}|=1$.
Suppose that $|A_{1}|>1$ and $|A_{2}|>1$. Given $\epsilon>0$ there are $\lambda_{1}>q_{B}(h_{A_{1}})-\epsilon/2$, $\lambda_{2}>q_{B}(h_{A_{2}})-\epsilon/2$, $x_{1},x_{2}\in\mathbb{R}^{k}$ such that
@@
We now have that $p$ is monotonic and sublinear and that $q_{B}$ is superlinear. Again we have that for any finite $A\subseteq\mathbb{R}^{k}$ and $\epsilon>0$ there is $\lambda$ such that $q_{B}(h_{A})-\epsilon<\lambda\leq q_{B}(h_{A})$ and $x\in\mathbb{R}^{k}$ with $\lambda B+x\subseteq\mathrm{conv}(A)$, and so
$p(h_{A})$ $\geq p(h_{\lambda B})=\lambda p(h_{B})>(q_{B}(h_{A})-\epsilon)\delta(B).$
@@
For $h_{A},h_{B}\in\mathcal{H}$ we write $h_{A}\preceq h_{B}$ if $h_{A}(x)\leq h_{B}(x)$ for all $x$, which, as $A$ and $B$ are convex bodies, holds exactly when $A\subseteq B$. The set $\mathcal{H}$ is a convex cone which, together with the partial order $\preceq$, satisfies the many axioms of a pre-ordered cone (see FuchssteinerLusky81convex for details). For each finite $B$ we have now satisfied the conditions for Theorem 1.2.5 of FuchssteinerLusky81convex :
@@
$q(h)=\delta(B)q_{B}(h)\leq\mu_{B}(h)\leq p(h).$
Since by definition $p(h_{\{a\}})=\delta(\{a\})=0$ for all $a\in\mathbb{R}^{k}$, we have $\mu_{B}(h_{\{a\}})=0$ for all $a\in\mathbb{R}^{k}$.
Now define $\delta_{B}$ by $\delta_{B}(A)=\mu_{B}(h_{A})$ for all finite $A$. Then $\delta_{B}$ vanishes on singletons, it is monotone, linear, and hence also sublinear. By Proposition 1, $(\mathbb{R}^{k},\delta_{B})$ is a linear semidiversity.
@@
$\delta_{B}(B)=\mu_{B}(h_{B})=\delta(B)q_{B}(h_{B})=\delta(B)$